OpenStudy (anonymous):

find limit: limit of x approaching -8 (x^2+3x-40)/(x+8)

OpenStudy (anonymous):

can you use the L'Hopital's rule here? It should be since the denom is 0 at the limit

OpenStudy (amistre64):

IF x=-8 has a finite value, it will cause the top to equal zero. in other words, if (x^2+3x-40) = (x+8)(....) then they both zero out and it has a hole at -8 that can be filled in

OpenStudy (anonymous):

or you can factorize the numerator and cancel the factor common to numerator and denominator. you will get the same answer

OpenStudy (amistre64):

x^2 +3x -40 0 -8x +40 -8 x -5 0 the good news is that it does factor, therefore to fill in the hole we can use the equivalent equation of: x-5

OpenStudy (anonymous):

@robtobey you copied the problem wrong. The denominator is 8+x and not 8x.

OpenStudy (anonymous):

\[\lim_{x \rightarrow -8} \frac{ x^2+3x-40 }{ x+8 }=\frac{ (x+8)(x-5) }{x+8}=x-5=-13 \]